# Video-MME: The First-Ever Comprehensive Evaluation Benchmark of Multi-modal LLMs in Video Analysis

### v2

@@ 3 Experiments
In this section, we evaluate a wide range of MLLMs on our Video-MME benchmark. We first introduce the evaluation settings, and then present the quantitative results for both open-source and closed-source models. Finally, we present case studies to provide an intuitive understanding, and investigate the effect of the modality information and duration length.
Table 3: The comparison of various benchmarks encompasses several key aspects: the total number of videos (#Videos), the number of clips (#Clips), the average duration of the videos (Len.), the number of QA pairs (#QA Pairs), the method of annotation (Anno., M/A means the manually/automatic manner), the average number of QA pair tokens (QA Tokens), the average number of subtitle tokens (Sub. Tokens), whether the videos cover multiple duration levels (Multi-level), whether the videos are sourced from a broad range of open domains (Open-domain), and whether provide subtitle together with audio information (Sub.&Aud.). Video-MME-S/M/L denotes the short/medium/long part. It is important to note that if a dataset includes multiple task formats, our comparison focuses solely on the multiple-choice segment.
@@ 3.1 Settings
We conduct the evaluation on 4 commercial models, i.e., GPT-4V, GPT-4o, Gemini 1.5 Flash, and Gemini 1.5 Pro. Representative open-source video MLLMs including Video-LLaVA, VideoChat2-Mistral, ST-LLM, Chat-UniVi-V1.5, LLaVA-NeXT-Video, and VILA-1.5 are evaluated as well. In addition, we also include advanced image MLLMs, i.e., Qwen-VL-Chat/Max and InternVL-Chat-V1.5, which usually can generalize to multi-image scenarios. We follow their official configurations and try to use more frames [^1] The numbers of the sampled frames are 10 for GPT-4V, 384 for GPT-4o, 8 for Video-LLaVA, 16 for VideoChat2-Mistral, 64 for ST-LLM, 64 for Chat-UniVi-V1.5, 32 for LLaVA-NeXT-Video, 8 for VILA-1.5, 4 for Qwen-VL-Chat/Max, and 10 for InternVL-Chat-V1.5. for evaluation. A special case is Gemini 1.5 Pro, because it supports extremely long multimodal contexts, so we take frames per second for both short and medium videos. For long videos, we capture a frame every two seconds to ensure the stability of the API. With respect to the setting of adding subtitles, all models except Gemini 1.5 Pro use the subtitles corresponding to the sampled video frames. For example, if you sample 10 frames, take the 10 subtitles that correspond to the time of those 10 frames. Gemini 1.5 Pro uses all subtitles due to the full video frame sampling. Besides, only Gemini 1.5 Pro supports the input of audios by now, whose results are listed in Table 5.
@@
The evaluation adopts the format of “whole video frames + whole subtitles/audios (optional) + question with prompt”. We try to use the model’s default prompt for multiple-choice questions, but if not we use a common prompt as:
@@
Considering the test sample in our benchmark is a multi-choice question with 4 options, we take accuracy as the evaluation metric, the random guess of which is 25%. The accuracy is calculated by matching the output of the model with the real one, without introducing any third party model such as ChatGPT. The example of accuracy calculation can be found on our project page.
3.2 Quantitative Results
Table 4: Performance of MLLMs on Video-MME with short, medium, and long durations, under the settings of “without subtitles” and “with subtitles”.
Models LLM Params     Short (%) Medium (%) Long (%) Overall (%)         w/o subs w/ subs w/o subs w/ subs w/o subs w/ subs w/o subs w/ subs Open & Closed-source Image MLLMs Qwen-VL-Chat [5]     7B     46.9 47.3 38.7 40.4 37.8 37.9 41.1 41.9 Qwen-VL-Max [5]     -     55.8 57.6 49.2 48.9 48.9 47.0 51.3 51.2 InternVL-Chat-V1.5 [9]     20B     60.2 61.7 46.4 49.1 45.6 46.6 50.7 52.4 Open-source Video MLLMs Video-LLaVA [30]     7B     45.3 46.1 38.0 40.7 36.2 38.1 39.9 41.6 ST-LLM [36]     7B     45.7 48.4 36.8 41.4 31.3 36.9 37.9 42.3 VideoChat2-Mistral [24]     7B     48.3 52.8 37.0 39.4 33.2 39.2 39.5 43.8 Chat-UniVi-V1.5 [19]     7B     45.7 51.2 40.3 44.6 35.8 41.8 40.6 45.9 LLaVA-NeXT-Video [74]     34B     61.7 65.1 50.1 52.2 44.3 47.2 52.0 54.9 VILA-1.5 [31]     34B     68.1 68.9 58.1 57.4 50.8 52.0 59.0 59.4 Closed-source MLLMs GPT-4V [48]     -     70.5 73.2 55.8 59.7 53.5 56.9 59.9 63.3 GPT-4o [49]     -     80.0 82.8 70.3 76.6 65.3 72.1 71.9 77.2 Gemini 1.5 Flash [54]     -     78.8 79.8 68.8 74.7 61.1 68.8 70.3 75.0 Gemini 1.5 Pro [54]     -     81.7 84.5 74.3 81.0 67.4 77.4 75.0 81.3
Table 5: Performance of Gemini 1.5 Pro on 6 major categories of Video-MME. The input modality includes frames only, frames with subtitles, and frames with audios.
Performance of Commercial Models.
As one of the pioneering commercial large models integrated with video comprehension capabilities, Gemini 1.5 Pro has achieved the best performance among its peers on Video-MME. As depicted in Table 4, with video frames as input alone, Gemini 1.5 Pro attains an accuracy of 75%, surpassing GPT-4V and GPT-4o by 15.1% and 3.1%, respectively. Table 5 shows the fine-grained performance of Gemini 1.5 Pro. Among the 6 major video categories, Gemini 1.5 Pro performs the best in Artistic Performance, achieving an accuracy of 78.8%, while performing the lowest in the Sports Competition category, with an accuracy of 68.6%. As video duration increases, Gemini 1.5 Pro’s performance declines (e.g., --6.7% from short to long videos), highlighting the model’s weakness in capturing long-range temporal relationships. Nevertheless, Gemini 1.5 Pro’s performance on long videos still surpasses almost all open-source models, except for VILA-1.5, on short videos, demonstrating its superior capabilities. In addition to visual frame input, Gemini 1.5 Pro’s support for additional modalities, including subtitles and audios, provides opportunities for further performance improvement. For example, Table 5 displays that using audios can increase accuracy by 6.2% for long videos, and the improvement in the multilingual category even reaches 16.7%. We can also see that the effect of subtitles and audios is different in these six categories. These motivate future research to develop versatile models that can support a wider range of modality inputs.
Performance of Open-sourced Models.
As shown in Table 4, among the 7B models, Chat-UniVi-V1.5 achieves the best performance with 40.6%. VILA-1.5 with 34B LLM achieves an accuracy of 59%, demonstrating its stronger capabilities, especially in the tasks of spatial reasoning, attribute perception, and information synopsis, as exhibited in Figure 3. Nevertheless, there still remains a significant gap between VILA-1.5 and Gemini Pro 1.5, particularly in counting problems, action recognition, and temporal perception, indicating substantial room for improvement. It is observed that adding subtitles can also help the open-sourced models. For example, the accuracy of LLaVA-NeXT-Video improves from 52% to 54.9%. It is regrettable that none of the open-sourced models support audio input.
Apart from video MLLMs, we also evaluate the performance of image MLLMs on Video-MME. Table 4 reveals that image-based Qwen-VL-Max and InterVL-Chat-V1.5 attain comparable performance to LLaVA-NeXT-Video, demonstrating their superior generalization capacity on sequential data, and the universality of Video-MME in both image and video MLLMs. Meanwhile, it also indicates that image understanding is the foundation of video understanding.
Figure 3: Performance on 12 types of tasks.
@@ 3.2 Quantitative Results
We conduct qualitative evaluation (using frames and subtitles) on the two cases in Figure 1. As analyzed in Section 2.2, these two cases comprehensively examine the model’s capabilities in OCR, attribute perception, object recognition, and long-range temporal reasoning, making them highly challenging. For the date-related question in Case 1, Video-LLaVA identifies the date (May 31st) from the frame at 01:10 and subtitles, but fails to perform reasoning based on context and incorrectly determines the year of the event, leading to the erroneous selection of option A. The remaining open-sourced models miscalculate the date 10 days after May 31st during the reasoning process, resulting in the incorrect choice of option C. For the event-related question in Case 2, Video-LLaVA, VideoChat2, and ST-LLM incorrectly associate the target person with nearby events, resulting in the selection of incorrect options A or C. In contrast, LLaVA-NeXT-Video and Gemini 1.5 Pro successfully identifies the events experienced by the target person throughout the video and demonstrates long-distance temporal modeling capabilities. They correctly link the target person’s injury at 03:35 with his reappearance at 27:30, identifying the true cause of the injury (option D). In summary, the questions in our benchmark pose significant challenges to the models, which motivates MLLMs to advance both their perception and reasoning capabilities.
3.3 Analysis
We conduct further analysis to explore the factors influencing the video understanding performance, e.g., additional information and video duration.
Could additional modalities benefit the performance?
Most of evaluations only take video frames as input, requiring models to answer questions solely based on visual contexts. However, many videos inherently contain extra information from other modalities, such as subtitles and audios. To understand their impact, we vary the combinations of input modalities in the evaluation, and report results in Table 5 and Figure 4. We can draw the following observations. (1) Introducing subtitles and audios can improve the results. For example, in the multilingual task shown in Table 5, with the addition of subtitles/audios, Gemini 1.5 Pro achieves +16.7%/+12.5% accuracy on long videos compared to the frame-only setting. This indicates that subtitles and audios provide some necessary information to answer the questions. (2) Subtitles and audios provide greater assistance in understanding long videos compared to short videos. For example, in Table 5, compared to only using frames, the addition of subtitles improves the model’s performance by 2.8% on short videos and by 10.1% on long videos. This is because test samples of long videos include more challenging reasoning questions, which requires the model to utilize subtitles and audio information for accurate responses. (3) For MLLMs, using subtitles is more effective than audios. Subtitles are usually transcriptions of audio, primarily capturing speech content, while audios encompass more ambient sounds. As can be seen from Table 5, whether it is short, medium, or long videos, subtitles bring higher improvement than audios. There are some exceptions in Table 5 and Figure 4, such as Multilingual, which may be due to the quality of the subtitles themselves. In addition, the audio also includes some content that subtitles can not express, such as singing and intonation.
Figure 4: The impact of incorporating additional modality inputs, i.e., subtitles and audios, on the performance of Gemini 1.5 Pro across different types of videos. We only show the results of 10 classes here, and those of the whole 30 classes can be found on our project page.
How MLLMs are robust to varied video duration?
In Table 4, we respectively compare the performance of different models on short, medium, and long videos. As video duration increases, both open-sourced and commercial models exhibit a significant decline in performance. For example, the accuracy of VILA-1.5 drops by 10% and 17.3% from short to medium and long videos, respectively, while Gemini 1.5 Pro’s accuracy decreases by 7.4% and 14.3%. There are three main reasons for the performance decline. (1) Increased proportion of difficult tasks. As shown in the bottom right corner of Figure 2, test samples for long videos contain a higher proportion of reasoning questions. These questions are more challenging than perception and recognition tasks, thus posing a greater challenge to the model’s capabilities. (2) Increased sparsity in frame sampling, leading to a reduction in effective input information. Ideally, for videos of varying lengths, models should sample video frames at a fixed fps to ensure consistent information density in frame sequences [56, 55]. However, many open-sourced models fix the number of input frames, e.g., 8 frames, resulting in excessively sparse information density as the video length increases. This sparsity prevents the model from retaining all useful visual semantics, hindering accurate predictions. Introducing additional modalities, e.g., subtitles, can effectively supplement the missing information [7]. (3) Increased difficulty in long context understanding. Although Gemini 1.5 Pro correspondingly increases the number of sampled frames in the long video, there is still a significant performance degradation. Understanding the long context of either single-modality (LLM) or multi-modality (MLLM) is always a great challenge.
4 Discussions
Our evaluation using Video-MME has revealed several critical insights into the current MLLMs and highlighted areas for future improvement. Here, we further discuss potential future directions.
Improving Long Context Modeling Capabilities of MLLMs.
One of the significant challenges identified in our evaluation is the decline in performance as video duration increases. For open-source models, the restricted input frames can become an information bottleneck for understanding the full content of long videos. Innovative approaches to context extension, both architectural and infrastructural, are essential. For instance, exploring techniques like ring attention [35], as investigated by large world models [34], and training-free context extension methods could be beneficial [2]. Additionally, developing architectures such as a temporal Q-Former to adaptively identify key frames in the video or compress video tokens to reduce computational overhead based on the questions posed is also worth exploring [56, 11]. In essence, improving long context modeling ability is crucial for the next generation of MLLMs to understand long sequential world dynamics effectively.
Building Datasets with Complex Temporal Understanding.
Our evaluation also highlights the demand for temporal reasoning oriented instruction-tuning datasets, considering that traditional video datasets with short video inputs, such as MSRVTT-QA and ActivityNet-QA. Although there have been efforts to construct high-quality datasets involving complex temporal reasoning over long videos [45, 23], the availability of such datasets is still insufficient compared to the text only [40, 46] and image datasets [27, 65]. The long-tailed nature of this data makes it challenging to acquire. Efforts towards better annotation paradigms such as human-in-the-loop frameworks [29] and automatic data synthesizing explorations are crucial [37]. Developing such datasets can better leverage advanced architectural innovations to provide MLLMs with sufficient training supervision for a robust understanding of the temporal dimension of videos.
5 Related Work
Advancements in MLLMs.
Recent advancements in MLLMs have seen notable progress [66, 14]. MLLMs typically comprise three core modules: (i) a vision encoder for visual feature extraction, (ii) a modality alignment module to integrate visual features into the embedding space of the language model, and (iii) an LLM backbone for decoding multi-modal context. CLIP [52] and SigLIP [70] are widely-used for image encoding, while LLaMA [58] and Vicuna [10] serve as popular choices for LLMs. The alignment module varies from simple linear projections [33, 75] to more complex architectures such as Q-Former [22, 11], and gated cross-attention layers substantiated by Flamingo and IDEFICS [1, 4]. Additionally, Fuyu-8B [6] introduces a novel framework mapping raw image pixels directly to the LLM embedding space. Regarding MLLMs for processing videos [23, 71, 43, 57, 32, 44, 19, 20, 64], the key difference lies in how they encode the video into vision tokens compatible with the LLMs. Representative work like Video-LLaMA [71] first uses a ViT [12] with an image Q-Former to encode individual frames and then employs a video Q-Former for temporal modeling. VideoChat2 [24] utilizes a video transformer to encode video features and subsequently implements a Q-Former [22] to compress video tokens. To empower video MLLMs with temporal localization capability [17, 51, 60], TimeChat [56] constructs time-sensitive instruction tuning datasets and encodes timestamp knowledge into visual tokens. VTimeLLM [16] proposes a LLaVA-like three-stage training method. However, the potential of MLLMs in processing sequential visual data is still under-explored. Therefore, we introduce Video-MME for full-spectrum, multi-modal evaluation of MLLMs in video analysis.
MLLM Benchmarks.
Alongside advancements in architecture, significant efforts have been made to improve benchmarking for MLLMs, guiding the development of the next generation of these models. Previous studies have integrated various aspects of evaluation, such as perception and cognitive capabilities, to create comprehensive benchmarks for assessing image MLLMs [13, 67, 38]. As image MLLMs have demonstrated exceptional performance in general perception tasks, benchmarks regarding scientific understanding [28], multi-modal mathematical reasoning [41, 73], and multi-disciplinary [69] capabilities have drawn increasing attention. For video MLLMs, similar efforts have been made to incorporate existing benchmarks [59, 26] for evaluating video understanding [24, 47]. Given the temporal nature of video modalities, specific benchmarks have been developed to address temporal understanding, highlighting the limitations of current video MLLMs in comprehending video content [29, 39].
In this work, we introduce a new high-quality video understanding benchmark, Video-MME. Compared to previous benchmarks, Video-MME includes a diverse set of videos of varying durations, supplemented with external modalities such as audios and subtitles. Additionally, it features human-annotated multi-level QA pairs, providing a comprehensive assessment framework for MLLMs. Our results indicate that open-source MLLMs still have a large gap with closed models. Our analysis and discussion further shed lights on the future development of MLLMs.
6 Conclusion
In this paper, we have introduced Video-MME, the first comprehensive multi-modal benchmark designed to evaluate MLLMs for video tasks. Our benchmark incorporates a diverse range of video types, varying temporal durations, and multiple data modalities, all annotated with high-quality, expert-labeled QA pairs. Our extensive evaluation of state-of-the-art MLLMs, including commercial and open-source models, highlights significant performance differences. Commercial models, particularly Gemini 1.5 Pro, demonstrate superior performance compared to open-source variants. The integration of subtitles and audio tracks significantly enhances video understanding, especially for longer videos. However, we observe a general decline in performance as video duration increases. These findings underscore the need for further advancements in handling longer multi-modal data. We hope Video-MME will inspire future research and development in improving the capabilities of MLLMs.
Acknowledgments
We appreciate the efforts made by Yu Bai, Fangyuan Liu, Yigeng Jiang, and Zezhong Wu in the construction of the benchmark.

### v3

@@ 4 Experiments
In this section, we evaluate a variety of MLLMs on our Video-MME benchmark. We begin by outlining the evaluation settings, followed by the quantitative results for both open-source and closed-source models Finally, we present case studies to provide an intuitive understanding, and investigate the effect of the modality and duration.
Figure 3: Comparison between four selected MLLMs on Video-MME across different tasks. The light blue in the outer circle stands for perception tasks and the orange represents reasoning tasks. As can be observed, counting problems are a joint bottleneck of current multimodal models in video understanding.
Table 5: Performance of Gemini 1.5 Pro across six major categories in the Video-MME benchmark. Evaluation includes three input modality settings: frames only, frames with subtitles, and frames with audio, highlighting the impact of multimodal inputs on model performance within each category. As one can observe, subtitles and audios contribute essential information for understanding the video.
@@ 4.1 Settings
We conduct the evaluation on 4 commercial models, i.e., GPT-4V [48], GPT-4o [49], Gemini 1.5 Flash [55], and Gemini 1.5 Pro [55]. Representative open-source video MLLMs including Video-LLaVA [33], VideoChat2-Mistral[27], ST-LLM [38], ShareGPT4Video [8],Chat-UniVi-V1.5 [23], LLaVA-NeXT-Video [73], VITA-1.0 [17], VITA-1.5 [18], and VILA-1.5 [34] are evaluated as well. In addition, we also include advanced image MLLMs, i.e., Qwen-VL-Chat/Max [5] and InternVL-Chat-V1.5 [11], which usually can generalize to multi-image scenarios. We follow their official configurations and try to use more frames for evaluation. The accuracy is computed by directly comparing the model’s output with the ground-truth answer, without utilizing any external models, such as ChatGPT.
4.2 Quantitative Results
Performance of Commercial Models.
As one of the pioneering commercial large models integrated with video comprehension capabilities, Gemini 1.5 Pro has achieved the best performance among its peers on Video-MME. As depicted in Table 4, with video frames as input alone, Gemini 1.5 Pro attains an accuracy of 75%, surpassing GPT-4V and GPT-4o by 15.1% and 3.1%, respectively. Table 5 shows the fine-grained performance of Gemini 1.5 Pro. Among the 6 major video categories, Gemini 1.5 Pro performs the best in Artistic Performance while performing the lowest in the Sports Competition category. As video duration increases, Gemini 1.5 Pro’s performance declines (e.g., --14.3% from short to long videos), highlighting the model’s weakness in capturing long-range temporal relationships. Nevertheless, Gemini 1.5 Pro’s performance on long videos still surpasses almost all open-source models, except for VILA-1.5, on short videos, demonstrating its superior capabilities. In addition to visual frame input, Gemini 1.5 Pro’s support for additional modalities, including subtitles and audios, provides opportunities for further performance improvement. For example, Table 5 displays that using audios can increase accuracy by 6.2% for long videos, and the improvement in the multilingual category even reaches 16.7%. We can also see that the effect of subtitles and audios is different in these six categories. These motivate future research to develop versatile models that can support a wider range of modality inputs.
Performance of Open-sourced Models.
As shown in Table 4, among the 7B models, VITA-1.5 achieves the best performance with 56.1%. VILA-1.5 with 34B LLM achieves an accuracy of 59%, demonstrating its stronger capabilities, especially in the tasks of spatial reasoning, attribute perception, and information synopsis (Figure 3). Nevertheless, there remains a significant gap between VILA-1.5 and Gemini Pro 1.5, particularly in counting problems, action recognition, and temporal perception, indicating substantial room for improvement. Regarding modality, adding subtitles consistently improves the performance of open-source models. Unfortunately, most of the open-source models do not support audio input, so audio evaluation is omitted.
Apart from video MLLMs, we also evaluate the performance of image MLLMs on Video-MME. Table 4 reveals that image-based Qwen-VL-Max and InterVL-Chat-V1.5 attain comparable performance to LLaVA-NeXT-Video, demonstrating their superior generalization capacity on sequential data, and the universality of Video-MME in both image and video MLLMs. It also indicates that image understanding is the foundation of video understanding.
4.3 Analysis
We conduct further analysis to explore the factors influencing the video understanding performance, e.g., additional modality information and video duration.
Could additional modalities benefit the performance?
Most evaluations use only video frames as input, requiring models to answer questions based solely on visual context. However, many videos inherently include additional information from other modalities, such as subtitles and audios. To understand their impact, we vary the combinations of input modalities in the evaluation. We can draw the following observations from the results in Table 5. (1) Introducing subtitles and audios can improve the results. For instance, in the multilingual task presented in Table 5, Gemini 1.5 Pro achieves a +16.7% and +12.5% accuracy improvement on long videos with the addition of subtitles and audios, respectively, compared to the frame-only setting. This suggests that subtitles and audios contribute essential information for answering the questions. (2) Subtitles and audios provide greater assistance in understanding long videos compared to short videos. For example, in Table 5, compared to only using frames, the addition of subtitles improves the model’s performance by 2.8% on short videos and by 10.1% on long videos. This is because test samples of long videos include more challenging reasoning questions, which require the model to utilize subtitles and audio information for accurate responses. (3) For MLLMs, using subtitles is more effective than audios. Subtitles are typically transcriptions of speech, focusing on verbal content, while audio includes additional ambient sounds. As shown in Table 5, subtitles consistently provide greater improvement than audio across different durations of videos. Multilingual can be regarded as an exception, possibly due to the quality of the subtitles.
How MLLMs are robust to varied video duration?
In Table 4, we respectively compare the performance of different models on short, medium, and long videos. As video duration increases, both open-sourced and commercial models exhibit a significant decline in performance. There are three main reasons for the performance decline. (1) Increased proportion of difficult tasks. As shown in the bottom right corner of Figure 2, test samples for long videos contain a higher proportion of reasoning questions, which poses a greater challenge to the model’s capabilities. (2) Increased sparsity in frame sampling, leading to a reduction in effective input information. Ideally, for videos of varying lengths, models should sample video frames at a fixed fps to ensure consistent information density in frame sequences [53, 52]. However, existing open-sourced models fix the number of input frames, e.g., 8 frames, resulting in excessively sparse information density as the video length increases. This sparsity prevents the model from retaining all useful visual semantics, hindering accurate predictions. Introducing additional modalities, e.g., subtitles, can effectively supplement the missing information [9]. (3) Increased difficulty in long context understanding. Although Gemini 1.5 Pro correspondingly increases the number of sampled frames in the long video, there is still a significant performance degradation. Understanding the long context of either single-modality (LLM) or multi-modality (MLLM) is always a great challenge.
5 Discussions
Our evaluation using Video-MME has revealed several critical insights into the current MLLMs and highlighted areas for future improvement. In this section, we provide more discussions on potential future directions.
Improving Long Context Modeling Capabilities of MLLMs.
One of the significant challenges identified in our evaluation is the decline in performance as video duration increases. For open-source models, the restricted input frames can become an information bottleneck for understanding the full content of long videos, which advocates for innovative approaches of both architectural and infrastructural context extension. For instance, exploring techniques like ring attention [37], as investigated by large world models [36], and training-free context extension methods could be beneficial [2]. Additionally, developing architectures such as a temporal Q-Former to adaptively identify key frames in the video or compress video tokens to reduce computational overhead based on the questions posed is also worth exploring [53, 13]. In essence, improving long context modeling ability is crucial for the next generation of MLLMs to understand long sequential world dynamics.
Building Datasets with Complex Temporal Understanding.
Our evaluation emphasizes the need for instruction-tuning datasets focused on temporal reasoning, particularly given the prevalence of traditional video datasets with short inputs, such as MSRVTT-QA and ActivityNet-QA. Although there have been efforts to construct high-quality datasets involving complex temporal reasoning in long videos [45, 26], the availability of such datasets remains limited compared to those for text [42, 46] and image [30, 65]. The long-tailed distribution of this data poses challenges for acquisition. Advancements in annotation methods, such as human-in-the-loop frameworks [32] and explorations into automatic data synthesis are crucial [39]. Developing these datasets will enable better use of architectural innovations, providing MLLMs with sufficient training supervision for robust temporal understanding in videos.
6 Conclusion
In this paper, we have introduced Video-MME, which is designed to evaluate MLLMs on video understanding tasks. Our benchmark incorporates a diverse range of video types, temporal durations, and data modalities with high-quality, expert-annotated QA pairs. Our extensive evaluation underscores the need for further advancements in handling longer multimodal data and we hope Video-MME will inspire future research and development in improving the capabilities of MLLMs.
Acknowledgments
This work was funded by National Natural Science Foundation of China under Grant 62441234. We appreciate the efforts made by Yu Bai, Fangyuan Liu, Yigeng Jiang, and Zezhong Wu in the construction of the benchmark.
Supplementary Material
7 Detailed Experimental Settings
Models.
We conduct a comprehensive evaluation on four commercial models and nine representative open-source video-based multimodal large language models. To further demonstrate the adaptability of our benchmark to multi-image scenarios, we also include three widely utilized image-based MLLMs as part of the evaluation. The complete list of models evaluated is provided below.
  • Commercial MLLMs: GPT-4V, GPT-4o, Gemini 1.5 Flash, and Gemini 1.5 Pro.
  • Open-source Video MLLMs: Video-LLaVA, ST-LLM, ShareGPT4Video, VideoChat2-Mistral, VILA-1.5, Chat-UniVi-V1.5, VITA-1.0, VITA-1.5, LLaVA-NeXT-Video.
  • Advanced Image MLLMs: Qwen-VL-Chat/Max and InternVL-Chat-V1.5.
Frame Extraction.
A standard approach involves extracting a sequence of frames from the video and interpreting the resulting multi-image inputs. For Gemini 1.5 Pro which supports extremely long multimodal contexts, we sample frames at 1 frame per second for short and medium videos, and at 1 frame every 2 seconds for long videos to ensure API stability. For all other models, frame extraction adheres to their respective official guidelines, uniformly sampling a specified number of frames from the video. The specific numbers of sampled frames are as follows: 10 frames for GPT-4V, 384 for GPT-4o, 8 for Video-LLaVA, 16 for VideoChat2-Mistral, 16 for ShareGPT4Video, 64 for ST-LLM, 64 for Chat-UniVi-V1.5, 32 for VITA-1.0, 16 for VITA-1.5, 32 for LLaVA-NeXT-Video, 8 for VILA-1.5, 4 for both Qwen-VL-Chat and Qwen-VL-Max, 10 for InternVL-Chat-V1.5.
Subtitle Utilization.
In the subtitle-enabled setting, all models utilize subtitles corresponding to the timestamps of the sampled video frames. For instance, if 10 frames are sampled from a video, the 10 subtitles that correspond to the respective timestamps of those frames are selected. This approach ensures accurate coherent synchronization between the visual and textual multimodal input for evaluation.
Evaluation.
@@ 7 Detailed Experimental Settings
The evaluation follows the format: “entire video frames + complete subtitles/audios (optional) + question with prompt.” Whenever possible, the model’s default prompt is utilized for multiple-choice questions. If unavailable, a standardized prompt is employed as follows:
@@
The accuracy is computed by extracting the model’s output using regular expressions and comparing it directly with the ground-truth answer, without relying on external judges like commonly-used ChatGPT.
8 Additional Analysis
How do MLLMs perform on the two highlighted cases in Figure 1?
@@ 8 Additional Analysis
We conduct qualitative evaluation (using frames and subtitles) on the two cases in Figure 1. As analyzed in Section 3.2, these two cases comprehensively examine the model’s capabilities in OCR, attribute perception, object recognition, and long-range temporal reasoning, making them highly challenging. For the date-related question in Case 1, Video-LLaVA identifies the date (May 31st) from the frame at 01:10 and subtitles, but fails to perform reasoning based on context and incorrectly determines the year of the event, leading to the erroneous selection of option A. The remaining open-sourced models miscalculate the date 10 days after May 31st during the reasoning process, resulting in the incorrect choice of option C. For the event-related question in Case 2, Video-LLaVA, VideoChat2, and ST-LLM incorrectly associate the target person with nearby events, resulting in the selection of incorrect options A or C. In contrast, LLaVA-NeXT-Video and Gemini 1.5 Pro accurately track the events involving the target individual across the entire video, showcasing robust long-range temporal modeling capabilities. They correctly link the target person’s injury at 03:35 with his reappearance at 27:30, identifying the true cause of the injury (option D). In summary, the questions in our benchmark pose significant challenges to the models, which motivates MLLMs to advance both their perception and reasoning capabilities.
Could additional modalities benefit the performance across categories?
Figure 4 presents the results of Gemini 1.5 Pro across the 30 subcategories of Video-MME, under the testing modes of frames, frames + subtitles, and frames + audio. The results indicate that subtitles and audio positively contribute to video understanding in multimodal large models. However, the extent of improvement provided by these modalities varies across different domains.
Figure 4: Evaluation results of Gemini 1.5 Pro across different video subcategories.